OpenStudy (anonymous):

Solve for x. 7(3 - x) = 8(4 - 2x) x = -53/23 x = -5/9 x = 11/23 x = 11/9

Parth (parthkohli):

=> 21 - 7x = 32 - 16x => -11 = -9x => 11 = 9x => x = 9/11

Parth (parthkohli):

Oops, x = 11/9

OpenStudy (anonymous):

7(3-x)=8(4-2x) 21-7x=32-16x -7x+16x=32-21 9x=11 x=11/9